# Pseudopotentials for Orbital-Free DFT: Capturing Nonlocality and Correcting Functional Approximants

### v1

@@ V.4 Compatibility of KEDF/LPP combinations
To evaluate the compatibility and performance of the proposed LPPs with several KEDF approximants, we examine the electron density of isolated atoms between two classes of KEDFs: semilocal (LKT[laricchia2011generalized], revAPBEK [constantin2011semiclassical]) and nonlocal (LMGP and revHC) KEDFs in comparison with densities computed with either TF0.2vW or TFvW employing the new LPPs, see Figure 5. To better understand the results for semilocal KEDFs, we consider the enhancement factor, $F(s)$, as a function of the reduced density gradient, $s({\mathbf{r}})$. Further details on the explicit forms of $F(s)$ for the different semilocal KEDFs discussed here (TFvW, TF0.2vW, LKT, and revAPBEK) are provided in the Supplementary Information section II. The left most panel of Figure 5 shows $F(s)$ where we notice that for small values of $s({\mathbf{r}})$, where the electron density is usually large LKT, revAPBEK , and TFvW have a noticeably higher $F(s)$ compared to TF0.2vW. The enhancement factor for TF0.2vW, due the small vW prefactor of 0.2, displays a slower upward behavior compared to the other functionals. Similarly for large values of $s({\mathbf{r}})$, which correspond to diffuse or tail regions of the density, the enhancement factor increases significantly for TFvW and LKT while it remains low for TF0.2vW and revAPBEK. The latter is due to its asymptotic behavior [laricchia2011generalized]. This analysis shows that LKT, revAPBEk and TFvW are likely more compatible with densities form TFvW in conjunction with PGBRV1.0 and PPSL1.0 than with densities from TF0.2vW in conjuction with PGBRV0.2 and PPSL0.2.
@@ VI Conclusions
In this study, we developed new sets of pseudopotentials for OF-DFT calculations targeting transition metal elements. Transition metals have historically been excluded from OF-DFT applications due to the challenges of constructing suitable local pseudopotentials. The core idea of this work is to adapt existing KS-DFT pseudopotentials (specifically, PSL and GBRV) by imposing that OF-DFT calculations with approximate noninteracting kinetic energy functionals reproduce, as closely as possible, the electron density from KS-DFT calculations of the element’s most stable condensed phase. This approach effectively performs a KS-to-OF inversion near the ion cores. This opens the possibility to investigate large-scale metallic systems and nanostructures at the OF cost (linear scaling with low prefactor). However, we point out that simulations based on OF-DFT for transition metals should still be considered “highly experimental” and not yet ready for broad use.

### v3

@@ V.4 Compatibility of KEDF/LPP combinations
To evaluate the compatibility and performance of the proposed LPPs with several KEDF approximants, we examine the electron density of isolated atoms between two classes of KEDFs: semilocal (TFvW, TF0.2vW, LKT[luo2020towards], revAPBEK[laricchia2011generalized]) and nonlocal (LMGP and revHC) KEDFs in comparison with densities computed with either TF0.2vW or TFvW employing the new LPPs, see Figure 5. To better understand the results for semilocal KEDFs, we consider the enhancement factor, $F(s)$, as a function of the reduced density gradient, $s({\mathbf{r}})$. Further details on the explicit forms of $F(s)$ for the various semilocal KEDFs (TFvW, TF0.2vW, LKT, and revAPBEK ) are provided in the Supplementary Information Section II. The left most panel of Figure 5 shows $F(s)$ where we notice that for small values of $s({\mathbf{r}})$, where the electron density is usually large LKT, revAPBEK , and TFvW have a noticeably higher $F(s)$ compared to TF0.2vW. The enhancement factor for TF0.2vW, due the small vW prefactor of 0.2, displays a slower upward behavior compared to the other functionals. Similarly for large values of $s({\mathbf{r}})$, which correspond to diffuse or tail regions of the density, the enhancement factor increases significantly for TFvW and LKT while it remains low for TF0.2vW and revAPBEK. The latter is due to its asymptotic behavior [laricchia2011generalized]. This analysis shows that LKT, revAPBEk and TFvW are likely more compatible with densities form TFvW in conjunction with PGBRV1.0 and PPSL1.0 than with densities from TF0.2vW in conjuction with PGBRV0.2 and PPSL0.2.
@@ VI Conclusions
In this study, we developed new sets of pseudopotentials for OF-DFT calculations targeting transition metal elements. Transition metals have historically been excluded from OF-DFT applications due to the challenges of constructing suitable local pseudopotentials. The core idea of this work is to adapt existing KS-DFT pseudopotentials (specifically, PSL and GBRV) by imposing that OF-DFT calculations with approximate noninteracting kinetic energy functionals reproduce, as closely as possible, the electron density from KS-DFT calculations of the element’s most stable condensed phase. This approach effectively performs a KS-to-OF inversion near the ion cores. Our work opens the possibility to investigate large-scale systems and nanostructures composed of transition metals at low cost (linear scaling with low prefactor). However, we also warn that simulations based on OF-DFT for transition metals should be considered highly experimental and not yet ready for broad adoption.